OpenStudy (anonymous):

If sin A = 4/5 and A is in Quadrant 2, and sin B =5/12 and B is in Quadrant 1, find the exact value of Sin (A-B)

OpenStudy (anonymous):

are you sure that sin B isn't 5/13? (5-12-13 triangle)

OpenStudy (anonymous):

im sure its not

OpenStudy (anonymous):

i think it must be otherwise we cant get an exact value

OpenStudy (anonymous):

ok my homework says 5/12 maybe thats a typo so lets assume its 5/13...what do we do

OpenStudy (anonymous):

ok then cos b = 12/13 , and cos a = -3/5 as its in 2nd quadrant so plud these values into the the identity sin (a-b) = sina cos b - cos a sin b = 4/5 * 12/13 -[ (-3/5 * 5/12) = 48/65 + 15/60

OpenStudy (anonymous):

= 48/65 + 1/4 = 192/260 + 65/260 = 257 / 260

OpenStudy (anonymous):

ok is that it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks very much

OpenStudy (anonymous):

i have one similar if you dont mind

OpenStudy (anonymous):

what if it was cos(A + B)

OpenStudy (anonymous):

ok cos (a + b) = cosa cosb - sina sinb = 12/13 * -3/5 - 5/13 * 4/5 = -36/65 - 20/65 = -56/65

OpenStudy (anonymous):

oh - theres an error in sin(a-b) - i used 5/12 instead of 5/13 so correct is sin (a-b) = 4/5 * 12/13 -[ (-3/5 * 5/13) = 48/65 + 15/65 = 63/65

OpenStudy (anonymous):

thanks very much..i have one last one...what if its tan(b/2)

OpenStudy (anonymous):

tan(b/2) - i cant remember the identity - ley me check it out - give me a few minutes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry jinnie i cant dig it out and i have to leave - theres is an identity linking tan (b/2) to sin b and cos b. you just need to plug in the values yu've got already

OpenStudy (anonymous):

ok thanks for the help

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i didnt have to be away for long and i remembered the formula tan(b/2) = sin b / (1 + cos b) = 5/13 / ( 1 + 12/13) = 5/13 / 25/13 = 5/13 * 13/25 = 1/5

OpenStudy (anonymous):

ok thanks very much

OpenStudy (anonymous):

these are all the questions i have pertaining to this concept...i still have other question but they are completely different concepts...would you be willing or do you have the time to help me more?